OpenStudy (anonymous):

How do I approach finding the volume of a solid generated by revolving the area bound by x1=2y-y^2, x2=0 about the x-axis?

OpenStudy (anonymous):

Plot the function

OpenStudy (anonymous):

yes, easy enough. will it have a differential in terms of x or y?

OpenStudy (anonymous):

x= 2y - 2y^2 x=0 is this right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=-y(y-2) x=0

OpenStudy (anonymous):